# Confined colloidal droplets dry to form circular mazes

### v1

@@ Appendix A Supplementary Figures
Fig. S1a(i) illustrates schematically (with a top-down and cross-sectional view) how the water slowly evaporates through the gap between the coverslip and the sticky tape that serves as spacer and confining walls of the flat sample cell containing 1 wtof 1.8 $\mu$m large TPM particles. Here, the water was fluorescently labelled with the green fluorescent dye calcein (excitation at 495 nm, emission at 515 nm) in order to visualize the water in the sample. Fluorescent images (512$\times$512 pixels) were taken with 50$\times$ magnification, while focused in the middle of the cell, and with a long exposure time to visualise where and how much of the fluorescent water is present. In all images, blue indicates a high water content (saturating the monitor for the green dye), red indicates a much lower water content and yellow an intermediate one.
@@
Fig. S1a(ii) shows a zoomed in image taken at the edge of the cell that demonstrates that water and air can be transported between the coverslip and the sticky tape. Fig. S1a(iii) is a 4$\times$4 grid of images taken to generate a picture of the entire sample at the moment finger formation starts. We refer to this drying period as stage 3 in the main text. This stage signifies the moment when the monolayer deposition on top and bottom surface stops and fingers start to form. The yellow to red coloring in the center of the droplet reflects the depression of the coverslip also depicted schematically in the sample cross-section in a(i) and b(i). This depression appears within the first 6 hours of evaporation and remains until the entire fingering pattern has formed. This is due to the capillary force the capillary bridge is developing [24] Assuming a wetting angle $\theta$ of roughly 45, an average height $h=60$$\mu$m and radius of r mm, we estimate the capillary force pushing the two surfaces together to be of the order of 15 N, which is sufficient to deform the coverslip while the sticky tape represents the counteracting force.
@@
Figure S1: Schematic and confocal images of a confined drying droplet containing 1.8 $\mu$m large TPM particles shown at different drying stages. a (i) Schematics of the top and cross-sectionional view of the sample at the early drying stage 1, showing the formation of air bubbles at the interface between the liquid and the sticky tape. (ii) Confocal image of the top surface of the cell that shows the fluorescent water trapped between the coverslip and sticky-tape spacer. Blue indicates a high and red a very low water content. Almost black regions bounded by the red to yellow inside the bubbles are dry and almost void of colloids. (iii) 4$\times$4 composite confocal image show the entire sample cell (total image size 1818$\times$1818 pixels), taken at the onset of stage 3. The scale bar is 1 mm. The schematics in b(i) illustrates the onset of monolayer deposition after the air bubbles have coalesced. The blue dashed rectangle corresponds to the composite confocal-image stacks in (ii), taken from the centre of the sample to the tap at the edge of the cell. The sharp black-to-green interface in the top, shows the outer air-suspension interface, while the integrated intensity decreases toward the centre of the sample, reflecting a dimple in the coverslip cause by capillary pressure. c Tiled confocal image of the sample, still focused on the bottom surface (after 4 hours and 20 minutes relative to a (iii)) when the entire pattern has formed, but is still wet. See the full drying sequence in Video V2.
@@
Fig. S1b(ii) is a tiling of confocal image stacks taken across the sample thickness (20 images taken over a sample thickness of 80 $\mu$m), shortly after the onset of finger formation. The entire strip is 7 mm long and 0.6 mm wide, starting from the middle of the sample (bottom) and ending at the tape at the sample edge (top). The drop of fluorescent intensity towards the center of the droplet (decreasing green levels) shows the depression of the sample due to capillary forces in the middle.
@@
Figure S2: Colloid distribution at the droplets’ air-water interface. a  Schematic of a droplet of the colloidal suspension placed in the middle of a sample cell, with both top-down and cross-sectional views. As this was placed in the middle rather than allowed to dry to this stage on its own, monolayer deposition has only just begun at the droplet edge at this stage. The blue dotted rectangle corresponds to the area shown in the two confocal images in b. The two images are individual slices from a stack on confocal images through the entire thickness of the sample (approx. 80 microns). The image on the left hand side is taken at z = 0, corresponding to the bottom surface of the cell, while the image on the left is taken at z = 65 microns (approx. 15 microns below the top surface). c  is an intensity plot of all the images of the interface stacked in the z direction from top to bottom. The inset plot shows the average bulk intensity from the bottom of the cell (z = 0) almost to the top of the cell at z = 65 microns.
@@
In Fig. S2a we illustrate schematically both a top and cross-sectional side view of a very small droplet of the colloidal suspension placed in the middle of the sample cell, such that it does not reach the cell’s sticky side wall. This allowed us to observe the interface between the vapour phase and the droplet confined between the top and bottom surface in a controlled way. Here we used 1.3 $\mu$m large TPM particles that were fluorescently labelled with BODIPY [21].
@@
Fig. S2b shows two confocal microscopy images taken at a 200 $\mu$m wide section across the air-droplet interface (blue dashed rectangle in the top view); they were recorded at the bottom ($z$ = 0) and near the top ($z$ = 65 $\mu$m) of the sample and are part of a $z$-stack of images leafing through the entire sample thickness, which is shown in Supporting Video V3. In Fig. S2c we plot the total intensity obtained for each $z$-layer of 2 $\mu$m thickness and averaged in $y$-direction that are shown in V3. The grey-scales in Fig. S2b indicate the intensity of the fluorescence coming form the colloids: Higher particle concentrations show stronger scattering intensity and thus a darker gray. The outside of the droplet is white as it is completely void of particles. At $z=0$$\mu$m, we see the onset of a colloidal monolayer deposition at around $x\approx 340$$\mu$m followed by a sharp intensity increase at $x\approx 310$$\mu$m (thick dark line) that coincides with the onset of the steep vapour-droplet interface. At even smaller $x$-values the scattering intensity levels off, reflecting a constant bulk-concentration of the fluorescently labelled colloids inside the droplet. From the intensity curves across the vapor-droplet interface we estimated the mean-width half-value (MWHV) as function of $z$. We found a MWHV of roughly 25 $\mu$m for all $z$-values. However, the colloid density of the particles at the bottom is higher than at the top.
@@
In Fig. S2c the intensity curves obtained from integrating the intensity (gray scale) in $xy$-pane are plotted as function of $z$ in a staggered way so that the intensity at the top-surface appears at lower intensity values and the bottom surface is shown at higher intensity values. In the inset we plot the intensity values as function of $z$ for each intensity curve, summed up from $x$ = 0 to 225 $\mu$m). The dashed line is an exponential fit to the intensity curve showing that near the air-water interface the colloids are Boltzmann distributed in $z$.
@@
Figure S3: Analysis of the drying kinetics shown in Fig. 3 in the main text. (a) Plot of the inscribed area $A_{1}$, which was estimated using method 1 (compact centre) and the area of the full pattern that includes also the area of the fingers. Snapshots of the respective thresholded patterns obtained for a given time at around $t=70$ hours are shown, where the area is yellow and the background is purple. The corresponding original photograph is also shown. (b) Average radius of the inscribed circle extracted from the areas of the inscribed circles plotted in (a) and using method 2, which correspond to the data for $r(t)$ shown in Fig. 3 in the main text. The corresponding fitting lines are plotted as solid lines and continued to the start of the drying process at $t=-166$ hours as dashed lines. Region 1 corresponds to the bubble formation and coalaescence, region 2 to the monolayer deposition and region 3 to the finger formation.
@@
In Fig. S3 we present different ways of analysing the shrinking area of the confined droplet as function of time, which is presented in Fig. 3 in the main text and video V4. Once finger deposition starts, the progressively shrinking drop area $A$, seen from top, can be described either by the total area of the fingers plus the inscribed circle or by the inscribed circle only. We developed a Python code to measure these two areas from the time-lapse video V4.
@@
In the analysis, we make use of two different methods of calculating the inscribed circle. Method A1 is based on applying a Sobel filter on the time-lapse images to find the lines of maximum contrast in the images. Then, we apply a Gaussian blur to be able to distinguish the centre from the fingers. We perform an image threshold based on Otsu’s algorithm to make the image binary, and then we perform image segmentation based on the Hoshen-Kopelman algorithm to separate all colloid-rich (i.e. whiter) areas of the images. Subsequently, we count the number of pixels in the central cluster, which is the total area of the inscribed circle $A_{1}$. We analyse the radius of this circle, which we denote $r_{1}(t)=r_{1}$.
@@
In the second method, we assessed the area $A_{2}(t)$ using the binarizing and thesholding methods of ImageJ. We estimated a slightly larger area than for method 1, but we could fit the resulting area $A_{2}(t)$ with the same functional form with a larger constant ($C=0.59$).
@@
In Fig. S4 we show the final drying patterns of several samples, each prepared under slightly different conditions. Typically, using lower concentrations of TPM particles lead to drying patters with a less dense monolayer deposition than the samples containing 1 wtTPM particles. Using double spacers with a cell height of $\sim 160\,\mu$m or higher particle concentrations lead to the onset of thick fingers that were kinetically hindered to form the transition to finer fingers at the centers. Adding the polymeric triblock surfactant Pluronic®  F127 suppressed finger formation completely.
Figure S4: Gallery of samples photographed in reflection mode.
Figure S5: Transition micrograph of a dried sample that was prepared with a suspension of 1 wtof 1.8 $\mu$m large TPM particles. The image was reconstructed from a 10 $\times$ 5 tiling of micrograpghs. The square cell is 24 $\times$ 12 mm large.

### v2

@@ Appendix A Supplementary Figures
Fig. S1a(i) illustrates schematically (with a top-down and cross-sectional view) how the water slowly evaporates through the gap between the coverslip and the double sided tape that serves as spacer and confining walls of the flat sample cell containing 1 wtof 1.8 $\mu$m large TPM particles. Here, the water was fluorescently labelled with the green fluorescent dye calcein (excitation at 495 nm, emission at 515 nm) in order to visualize the water in the sample. Fluorescent images (512$\times$512 pixels, with pixel size 4.55 $\mu$m) were taken with a 5$\times$ objective lens, while focused in the middle of the cell (approximately 40 $\mu$m from the bottom surface), and with a long exposure time to visualise where, and how much, of the fluorescent water is present. In all images, blue indicates a high water content (saturating the monitor for the green dye), red indicates a much lower water content and yellow an intermediate one.
@@
Fig. S1a(ii) shows a zoomed in image taken at the edge of the cell that demonstrates that water and air can be transported between the coverslip and the double sided tape. Fig. S1a(iii) is a 5$\times$5 grid of images taken, as described above, to generate a picture of the entire sample at the moment finger formation starts. We refer to this drying period as stage 3 in the main text. This stage signifies the moment when the monolayer deposition on top and bottom surface stops and fingers start to form. The yellow to red coloring in the center of the droplet reflects the depression of the coverslip also depicted schematically in the sample cross-section in a(i) and b(i). This depression appears within the first 6 hours of evaporation and remains until the entire fingering pattern has formed. This is due to the capillary force produced by the capillary bridge [24]. Assuming a wetting angle $\theta$ of roughly 45, an average height $h=60$$\mu$m and radius of $r=5$mm, we estimate the capillary force pushing the two surfaces together to be of the order of 0.15 N, which is sufficient to deform the coverslip while the double sided tape represents the counteracting force.
@@
Figure S1: Schematic and confocal images of a confined drying droplet containing 1.8 $\mu$m large TPM particles, and fluorescently dyed water, shown at different drying stages. a (i) Schematics of the top-down and cross-sectional view of the sample at stage 1 of the drying process, showing the formation of air bubbles at the interface between the liquid and the double sided tape. (ii) Confocal image of the top surface of the cell that shows the fluorescent water trapped between the coverslip and double sided tape spacer. Blue indicates a high, and red a very low, water content. The almost black regions bounded by the red to yellow inside the bubbles are dry and almost void of colloids. (iii) 5$\times$5 tiled confocal image show the entire sample cell (total image size 2560$\times$2560 pixels), taken at the onset of stage 3. The scale bar is 1 mm. The schematics in b(i) illustrates the onset of monolayer deposition after the air bubbles have coalesced. The blue dashed rectangle corresponds to the composite confocal-image stacks in (ii), taken from the centre of the sample to the tape at the edge of the cell. The sharp black-to-green interface at the top, shows the outer air-suspension interface, while the integrated intensity decreases toward the centre of the sample, reflecting a dimple in the coverslip cause by capillary pressure. c A tiled confocal image of the sample (after 4 hours and 20 minutes relative to a (iii)) when the entire pattern has formed, but is still wet. See the full drying sequence in Video V2.
@@
Fig. S1b(ii) is a tiling of confocal image stacks taken across the sample thickness (50 images taken over a sample thickness of 80 $\mu$m), shortly before the onset of finger formation. The entire strip is 7 mm long and 0.6 mm wide, starting from the middle of the sample (bottom) and ending at the tape at the sample edge (top). The drop of fluorescent intensity towards the center of the droplet (decreasing green levels) shows the depression of the sample due to capillary forces in the middle.
@@
Figure S2: Colloid distribution at the droplets’ air-water interface. a  Schematic of a droplet of the colloidal suspension placed in the middle of a sample cell, with both top-down and cross-sectional views. As this was placed in the middle, rather than allowed to dry to this stage on its own, monolayer deposition has only just begun at the droplet edge at this stage. The blue dotted rectangle corresponds to the area shown in the two confocal images in b. The two images are individual slices from a stack on confocal images through the entire thickness of the sample (approx. 80 $\mu$m). The image on the left hand side is taken at z = 0, corresponding to the bottom surface of the cell, while the image on the left is taken at z = 65 microns (approx. 15 $\mu$m below the top surface). c  An intensity plot of all the images within the confocal image stack of the interface, stacked in the z direction from top to bottom. The inset plot shows the average bulk intensity from the bottom of the cell (z = 0) almost to the top of the cell at z = 65 $\mu$m.
@@
In Fig. S2a we illustrate schematically both a top-down and cross-sectional view of a small droplet of the colloidal suspension placed in the middle of the sample cell, such that it does not reach the wall of the cell made by the double sided tape. This allowed us to observe the interface between the vapor phase and the droplet confined between the top and bottom surface in a controlled way. Here we used 1.3 $\mu$m large TPM particles that were fluorescently labelled with BODIPY [21].
@@
Fig. S2b shows two confocal microscopy images taken across a 200 $\mu$m wide section at the air-droplet interface (blue dashed rectangle in the top view); they were recorded at the bottom ($z$ = 0) and near the top ($z$ = 65 $\mu$m) of the sample and are part of a $z$-stack of images leafing through the entire sample thickness, which is shown in Supporting Video V3. In Fig. S2c we plot the total intensity obtained for each $z$-layer of 2 $\mu$m thickness and averaged in $y$-direction that are shown in V3. The grey-scales in Fig. S2b indicate the intensity of the fluorescence coming form the colloids: higher particle concentrations show stronger scattering intensity and thus a darker gray. The outside of the droplet is white as it is completely void of particles. At $z=0$$\mu$m, we see the onset of a colloidal monolayer deposition at around $x\approx 340$$\mu$m followed by a sharp intensity increase at $x\approx 310$$\mu$m (thick dark line) that coincides with the onset of the steep vapour-droplet interface. At even smaller $x$-values the scattering intensity levels off, reflecting a constant bulk concentration of the fluorescently labelled colloids inside the droplet. From the intensity curves across the vapor-droplet interface we estimated the mean-width half-value (MWHV) as function of $z$. We found a MWHV of roughly 25 $\mu$m for all $z$-values. However, the colloid density of the particles at the bottom is higher than at the top.
@@
In Fig. S2c the intensity curves obtained from integrating the intensity (gray scale) in the $xy$-plane are plotted as function of $z$ in a staggered way so that the intensity at the top-surface appears at lower intensity values and the bottom surface is shown at higher intensity values. In the inset we plot the intensity values as function of $z$ for each intensity curve, summed up from $x$ = 0 to 225 $\mu$m). The dashed line is an exponential fit to the intensity curve showing that near the air-water interface the colloids are Boltzmann distributed in $z$.
@@
Figure S3: Analysis of the drying kinetics shown in Fig. 3 in the main text. a Plot of the inscribed area $A_{1}$, which was estimated using method 1 (compact centre) and the area of the full pattern that includes also the area of the fingers. Snapshots of the respective thresholded patterns obtained for a given time at around $t=70$ hours are shown, where the area is yellow and the background is purple. The corresponding original photograph is also shown. b Average radius of the inscribed circle extracted from the areas of the inscribed circles plotted in a and using method 2, which correspond to the data for $r(t)$ shown in Fig. 3 in the main text. The corresponding fitting lines are plotted as solid lines and continued to the start of the drying process at $t=-166$ hours as dashed lines. Region 1 corresponds to the bubble formation and coalaescence, region 2 to the monolayer deposition and region 3 to the finger formation.
@@
In Fig. S3 we present different ways of analysing the shrinking area of the confined droplet as function of time, which is presented in Fig. 3 in the main text and video V4. Once finger deposition starts, the progressively shrinking drop area $A$, as seen from the top, can be described either by the total area of the fingers plus the inscribed circle or by the inscribed circle only. We developed a Python code to measure these two areas from the time-lapse video V4.
@@
In the analysis, we make use of two different methods of calculating the inscribed circle. Method A1 is based on applying a Sobel filter on the time-lapse images to find the lines of maximum contrast in the images. Then, we apply a Gaussian blur to be able to distinguish the centre from the fingers. We perform an image threshold based on Otsu’s algorithm to binarise the images, and then we perform image segmentation based on the Hoshen-Kopelman algorithm to separate all colloid-rich (i.e. whiter) areas of the images. Subsequently, we count the number of pixels in the central cluster, which is the total area of the inscribed circle $A_{1}$. We analyse the radius of this circle, which we denote $r_{1}(t)=r_{1}$.
@@
In the second method, we assessed the area $A_{2}(t)$ using the WebPlotDigitizer from Automeris LLC. In this method we estimated a slightly larger area than for method 1, as it considered the outer perimeter of the inscribed circle but excluded the finger. We could fit the resulting area $A_{2}(t)$ with the same functional form with a larger constant ($C=0.59$).
@@
In the first row of Fig. S4, we show the final drying patterns of several other samples, made in the same way as those shown in the main text, to demonstrate the highly reproducible nature of these patterns. In the second row, we then show 4 more samples, each prepared under slightly different conditions. Typically, using lower concentrations of TPM particles lead to similar drying patterns, but with a less dense monolayer deposition than the samples containing 1 wtTPM particles. In contrast, using higher concentrations of TPM particles lead to similar monolayer deposition behaviour as for the samples containing 1 wtTPM particles, but produced much less clearly defined patterns in the centre. Using a double spacer with a cell height of $\sim 160\,\mu$m lead to the onset of thick fingers that were kinetically hindered to form the transition to finer fingers at the centers, but we still observe monolayer deposition as before. Adding the polymeric triblock surfactant Pluronic®  F127 suppressed finger formation completely. In Fig. S5, we also show that we observed the same pattern formation phenomenon with a rectangular geometry.
Figure S4: Gallery of samples photographed in reflection. The top row shows 4 other samples with the same composition as those in the main text, showing high levels of reproducibility. The bottom row shows some other samples, each with a different modification. From left to right: 0.5wtconcentration of TPM, 2wtconcentration of TPM, double-spaced surfaces (cell height of 160 $\mu$m), and 1.3 $\mu$m large TPM colloids that have been surface functionalised with Pluronic®  F127.
Figure S5: Transition micrograph of a dried sample that was prepared with a suspension of 1 wtof 1.8 $\mu$m large TPM particles. The image was reconstructed from a 12$\times$6 tiling of micrograpghs. The square cell is 24 $\times$ 12 mm large.